A floodlight at ground level 6 feet from a bush 4 feet tall casts a shadow of the bush on a building that is 24 feet from the bush. What is the height of the shadow?
Answer by nyc_function(2741)

Let h = height of shadow


6/24 = 4/h


24(4) = 6h

96 = 6h

96/6 = h

16 feet = h


Done!